| Date | Event |
| May 2010 | MWR(HATPRO and 90/150) installed at Summit |
| 4 Jun 2010 | LN2 calibration |
| 15 Jun - 11 Jul 2010 | Data collected but tweaked many times (old config) |
| 12 jul 2010 | Start calibrated file production |
| 5 oct 2010 | Possible mirror position change (seen in 52 and 54 GHz) |
| 23-25 Oct 2010 | Internal temps unstable on HATPRO, unavailiable for 90/150 and likely unstable |
| 24 oct 2010 | Tip curve difference off on all channels of HATPRO |
| 27 Oct - 1 Nov 2010 | Internal temps unstable on HATPRO, plugged vents thereafter |
| 29 Oct 2010 | Gain calibration taken 3x as often (SUMMIT_VERSION_09_SCAN.MBF) |
| 2 Nov 2010 | Thermal blankets around HATPRO, restarted |
| 3 Nov 2010 | LN2 calibration. Bad cal because of incorrect mirror position |
| 17 Nov 2010 | HATPRO taken off the roof for repairs |
| 17 Nov 2010 - 19 Jan 2011 | No "*SL.HKD.NC files." Housekeeping files go with 90/150 |
| 30 Dec 2010 | HATPRO mirror assembly repaired and back on the roof |
| 12-19 Jan 2011 | HATPRO error condition, no data, very cold temps |
| 15-20 Mar 2011 | HATPRO and 90/150 error condition-connectivity problems, sparse data, very cold temps |
| 28-30 April 2011 | The MSF was moved, HATPRO and 90/150 were powered down |
| 12 June 2011 | Performed LN2 calibration on both the HATPRO and HF |
| 10-23 June 2011 | Connectivity issues with the HF MWR |
| 23 June 2011 | replaced fiber optic cable on HF with the spare copper cable |
| 22 July 2011 | tested leveling the MSF ~1430UTC |
| 25 July 2011 | leveled the MSF ~1630UTC. Raised the NE corner |
| 16 Aug 2011 | Performed Ln2 calibration on both HATPRO and HF |
| 16 Aug 2011 | rotated both the HATPRO and the HF by a few degrees since the Swiss Tower might have been in the FOV since the move in the Spring |
| 16 Aug 2011 | Added a foam insulation piece between the stand and the MWRs with additional aerogel insulation over the MWRs |
| 23 April 2012 | Performed Ln2 calibration on both HATPRO and HF in addition to the "flop" scan |
| 10 May 2012 | MSF powered down. The MWR insulation and tape around the power switches was removed and then replaced. |
| 17 May 2012 | MSF moved. MWRs powered down. Roof shoveled clean. |
| 19 May 2012 | MWRs are rotated in order to align the FOV with each other. |
| 10-18 July 2012 | Unstable internal temps in the HATPRO MWR (too warm). |
| 12 July 2012 | Implemented the calibration values from the 16 August 2012 LN2 calibration. |
| 13 July 2012 | Removed insulation from HATPRO MWR.Internal temps still too warm |
| 17 July 2012 | Removed tape from vent covers on HATPRO and attached vent covers. |
| 21 July 2012 | leveling work on the MSF ~1700UTC. |
| 25 July 2012 | leveled the MSF ~1745UTC. |
| 3 Aug 2012 | MWRs leveled (HATPRO and HF) |
| 11-13 Aug 2012 | Unstable internal temps in the HATPRO MWR. |
| 13 Aug 2012 | Re-applied insulation over MWRS and plugged vents: HATPRO(~1800 UTC) and HF(1700 UTC) |
| 15 Aug 2012 | LN2 calibration 1 on MWRs. HATPRO flop cal (1216-1234 UTC) and HF(1308-1330 UTC) |
| 15 Aug 2012 | LN2 calibration 2 on MWR. HATPRO flop cal (1851-1939 UTC) |
| 16 Aug 2012 | LN2 calibration 3 on MWR. HATPRO flop cal (1254-1341 UTC) |
| 25 April 2013 | LN2 calibration 1 on both MWRs. HATPRO flop cal |
| 26-28 April 2013 | Additional LN2 calibrations on HATPRO MWR. HATPRO flop cal |
| 8-14 May 2013 | Additional LN2 calibrations on HATPRO MWR. HATPRO flop cal |
| 21 April 2014 | HATPRO removed from roof of MSF for repairs/refurbishment |
| 22 April 2014 | No "*SL.HKD.NC files." Housekeeping files go with 90/150 |
| 15 May 2014 | MSF moved. MWRs powered down during move. |
| 19 Aug 2014 | HATPRO reinstalled on the roof after repairs and refurbishment |
| 19 Dec 2014 | leveled the MSF |
| 23 Jan 2015 | adjusted the level of the short axis of the HATPRO and HF MWRs |
| 26 Dec 2015 | Red flag on the noise diode of the HATPRO. Data compromised. |
| 11 Jan 2016 | Red flag on the noise diode of the HATPRO. HATPRO removed for repairs in Germany. |